Grade 11General Physics

The sum of magnitudes of two forces acting at a point is 18 and the magnitude of their resultant is 12 if the resultant is at 90 degree with the force of smaller magnitude what are the magnitudes of Forces

Susmita
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
Suppose that A and B be the powers where A is more noteworthy than B , acting at a point P slanted edge . 
Here A+B = 18 ....(1) and the resultant R =12 N. Develop a parallelogram with sides with PC = A, PN = B, with point CPN as an insensitive edge and PC along the x axis.Draw the corner to corner PK to the parellelogram OCKN so that edge CPN =90 degree .The inclining PN = resultant R which is opposite to the s maller drive Q . consider the triangle CPN , PC = A, CK = PN = B, and edge CPN =90degree.That is triangle CPN is a privilege calculated triangle with CN = P as hypotenuse. At that point A^ 2 = B^2 +R^2 or A^2 - B^2 = R^2 or (A+B) (A - B) = R^2 or A - B = R^2/A+B that is A-B = 144/18 = 8 ...(2) Adding (1) and (2) 2A = 26 or A = 13 N . At that point B = 5 N.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)